1 The Lower Sorbian language

The file lsorbian.dtx^{1} It defines all the language-specific macros for Lower Sorbian.

The macro \LdfInit takes care of preventing that this file is loaded more than once, checking the category code of the @ sign, etc.

1 ⟨∗code⟩

 2 \LdfInit{lsorbian}\captionslsorbian

When this file is read as an option, i.e. by the \usepackage command, lsorbian will be an ‘unknown’ languagein which case we have to make it known. So we check for the existence of \l@lsorbian to see whether we have to do something here.

 3 \ifx\l@lsorbian\@undefined
 4 \@nopatterns{Lsorbian}
 5 \adddialect\l@lsorbian\l@usorbian\fi

The next step consists of defining commands to switch to (and from) the Lower Sorbian language.

\captionslsorbian The macro \captionslsorbian defines all strings used in the four standard documentclasses provided with L^{A}TEX.

 6 \addto\captionslsorbian{%
 7 \def\prefacename{Zawod}%
 8 \def\refname{Referency}%
 9 \def\abstractname{Abstrakt}%
10 \def\bibname{Literatura}%
11 \def\chaptername{Kapitl}%
12 \def\appendixname{Dodawki}%
13 \def\contentsname{Wop\'simje\'se}%
14 \def\listfigurename{Zapis wobrazow}%
15 \def\listtablename{Zapis tabulkow}%
16 \def\indexname{Indeks}%
17 \def\figurename{Wobraz}%
18 \def\tablename{Tabulka}%
19 \def\partname{\'Z\v el}%
20 \def\enclname{P\'si\l oga}%
21 \def\ccname{CC}%
22 \def\headtoname{Komu}%
23 \def\pagename{Strona}%
24 \def\seename{gl.}%
25 \def\alsoname{gl.~teke}%
26 \def\proofname{Proof}%  <-- needs translation
27 }%

\newdatelsorbian The macro \newdatelsorbian redefines the command \today to produce Lower Sorbian dates.

28 \def\newdatelsorbian{%
29 \def\today{\number\day.~\ifcase\month\or
30 januara\or februara\or m\v erca\or apryla\or maja\or junija\or
31 julija\or awgusta\or septembra\or oktobra\or

1The file described in this section has version number v1.0d and was last revised on 1996/12/23. It was written by Eduard Werner (edi@kaihh.hanse.de).

32 nowembra\or decembra\fi
33  \space \number\year}}

\olddatelsorbian The macro \olddatelsorbian redefines the command \today to produce old-style Lower Sorbian dates.

34 \def\olddatelsorbian{%
35 \def\today{\number\day.~\ifcase\month\or
36  wjelikego ro\v zka\or
37  ma\l ego ro\v zka\or
38  nal\v etnika\or
39  jat\v sownika\or
40  ro\v zownika\or
41  sma\v znika\or
42  pra\v znika\or
43  \v znje\'nca\or
44  po\v znje\'nca\or
45  winowca\or
46  nazymnika\or
47  godownika\fi \space \number\year}}

The default will be the new-style dates.

48 \let\datelsorbian\newdatelsorbian

\extraslsorbian \noextraslsorbian The macro \extraslsorbian will perform all the extra definitions needed for the lsorbian language. The macro \noextraslsorbian is used to cancel the actions of \extraslsorbian. For the moment these macros are empty but they are defined for compatibility with the other language definition files.

49 \addto\extraslsorbian{}
50 \addto\noextraslsorbian{}

The macro \ldf@finish takes care of looking for a configuration file, setting the main language to be switched on at \begin{document} and resetting the category code of @ to its original value.

51 \ldf@finish{lsorbian}

52 ⟨/code⟩